I don't know how to set this problem up: Two cyclists start at the same point and travel in opposite directions. One travels 3mph faster than the other. In 2 hours they are 402 miles apart. How fast is each traveling?
Answer by JulietG(1812)

The total distance is 402 miles.
Let the faster cyclist be A and the slower, B.
How many miles traveled in one hour? 201
A + B = 201
A = B + 3
Substitute the value of A from the second equation into the first.
(B+3) + B = 201
2B + 3 = 201
Subtract 3 from each side
2B = 198
B = 99
If B = 99 mph, then A = 102 mph
.
Obviously impossible ;-), unless they are MOTORbikes (doubtful even then!)